Please help, I am having problems with this one, thanks
Media services charges $40 for a phone and $20/month for its economy plan. The equation C= 20t + 40 describes the total cost, C, of operating a media service phone for t months. the total cost for 4 months of service is $____.
If a customer has only $90 available, how many months of service can she receive? Would it be 3, 2 1/3, 2 1/2 or 2?

Found 2 solutions by Theo, nerdybill:
Answer by Theo(13342)

C = 20t + 40
4 months service would be 20*4 + 40 = 80 + 40 = 120
-----
if C = 90, then equation becomes 90 = 20t + 40
subtract 40 from both sides of this equation to get 90 - 40 = 20t which becomes:
50 = 20t
divide both sides of this equation by 20 to get 50/20 = t which means that t = 50/20 = 5/2 = 2 and 1/2 months of service.
-----
your answer will depend on whether the service allows partial months.
-----
if it does, then your answer is 2 and 1/2 months.
you will pay 40 for the phone plus 40 dollars for 2 months plus 10 dollars for 1/2 a month to make a total of 40 + 40 + 10 = 90.
-----
if it doesn't allow partial months, then the best you can do is 2 months of service and you get to keep the extra 10.
2 * 20 = 40 + 40 for the phone = 80 leaving you 10 dollars extra.
-----
since they did not tell you that the service did not allow partial months then you should assume that they would allow it and your answer would be 2 and 1/2 months for 90.
-----

Answer by nerdybill(7384)

Media services charges $40 for a phone and $20/month for its economy plan. The equation C= 20t + 40 describes the total cost, C, of operating a media service phone for t months. the total cost for 4 months of service is $____.
If a customer has only $90 available, how many months of service can she receive? Would it be 3, 2 1/3, 2 1/2 or 2?
.
The problem gives you the "cost" as a function of 't' (months):
C(t) = 20t + 40
where t is the number of months
.
For 4 months then:
C(t) = 20t + 40
C(4) = 20(4) + 40 = 80 + 40 = $120
the total cost for 4 months of service is $120
.
If a customer has only $90 available:
C(t) = 20t + 40
90 = 20t + 40
50 = 20t
50/20 = t
5/2 = t
or
2 1/2 months = t